OpenStudy (anonymous):

Can someone please solve this for me? 6x^2 – 24x = 126.

OpenStudy (anonymous):

6x^2 - 24x - 126 = 0 Factor out 6 x^2 - 4x - 21 = 0/6 = 0 factors of 21 are 3 and 7, difference of 4 = can factor (x-7)(x-3)=0 x=7 or x=3

OpenStudy (anonymous):

oops i missed a sign

OpenStudy (anonymous):

(x-7)(x+3)=0 x = 7 or x= -3

OpenStudy (anonymous):

x=-1.567 x=2.234

jhonyy9 (jhonyy9):

6x2 -24x-126=0 divide both sides by 6 x2 -4x -21=0 this can be factored that -4= -7+3 and -21= (-7)*(3) than will be (x-7)(x+3)=0 so x-7=0 so x=7 and x+3=0 so x= -3